OpenStudy (anonymous):

Can someone help me solve this equation? n/7-5n/14=1/7

OpenStudy (dangerousjesse):

Sure, let me type out the steps :)

OpenStudy (anonymous):

Thank you soo much!

OpenStudy (dangerousjesse):

Multiply both sides by a constant to simplify the equation. Multiply both sides of \(\frac{-3 n}{14} = \frac{1}{7}\) by \(\frac{-14}{3:}\) \(\frac{-14 (-3) n}{3}\div14 = \frac{-14}{3}\div 7\) Express ((-14)/3)/7 as a single fraction. \(\frac{-14}{3}\div 7 = \frac{-14}{3×7}\): \(\frac{-14 (-3) n}{3×14} = \frac{-14}{3×7}\) In \(\frac{-14 (-3) n}{3×14}\), divide -14 in the numerator by 14 in the denominator. \(\frac{-14}{14} = \frac{14 (-1)}{14} = -1:\) \(\frac{-3-1 n}{3 = \frac{-14}{3×7}\) In -(-3 n)/3, divide -3 in the numerator by 3 in the denominator. \(\frac{-3}{3} = \frac{3 (-1)}{3} = -1:\) -\([-1] n = \frac{-14}{3×7}\) Multiply all instances of -1 in \(-(-n).\) \(-1^2 = 1:\) \(n = \frac{-14}{3×7}\) In \(\frac{-14}{3×7}\), divide -14 in the numerator by 7 in the denominator. \(\frac{-14}{7} = \frac{7 (-2)}{7} = -2:\)

OpenStudy (dangerousjesse):

Gah, I hate when OS screws with the fractions. Sorry if it looks weird.